OpenStudy (anonymous):

g = 3h - 4i solve for h

OpenStudy (saifoo.khan):

g + 4i = 3h

OpenStudy (saifoo.khan):

g + 4i = 3h divide 3 from both sides, (g+4i)/3 = h